The curve y = 1 -(1/4)x^2 intersects the positive side of the x-axis at A and the y-axis at B. O is the origin. Calculate the volume generated when the finite area bounded by BO, OA, and the arc AB is rotated through four right angles
i. about the x-axis
ii. about the y-axis.
Give each answer as a multiple of π.
1
Expert's answer
2021-08-09T16:39:44-0400

FORMULAS: "V=\displaystyle\int A dx", or respectively "\displaystyle\int Ady" where "A" stands for the area of the typical disc. Another words: "A = \pi r^2" and "r=f(x)" or "r=f(y)" depending on the axis of revolution.


  • The volume of the solid generated by a region under "f(x)" bounded by the "x" - axis and vertical lines "x=a" and "x=b", which is revolved about the x-axis is
"V=\pi \displaystyle\int\limits_a^by^2dx=\pi\displaystyle\int\limits_a^b\Big[f(x)\Big]^2dx"


  • The volume of the solid generated by a region under "f(y)" (to the left of "f(y)" bounded by the "y" - axis, and horizontal lines "y=c" and "y=d" which is revolved about the y-axis.
"V=\pi \displaystyle\int\limits_c^dx^2dy=\pi\displaystyle\int\limits_c^d\Big[f(y)\Big]^2dy"




SOLUTION: "y=1-\dfrac{x^2}{4}, ~~~x=\pm2\sqrt{1-y}~;"


"a)~ V_x = 2 \pi \displaystyle\int_{0}^{2} \Big[f(x)\Big]^2 dx = 2\pi \displaystyle\int_{0}^{2}\Bigg[\dfrac{4-x^2}{4}\Bigg]^2dx=2\pi \cdot \dfrac{16}{15}=\dfrac{32}{15}\pi;""b)~ V_y=2\pi\displaystyle\int_{0}^{1}\Big[f(y)\Big]^2dy=2\pi\displaystyle\int_{0}^{1}\Bigg[2\sqrt{1-y}\Bigg]^2dy=2\pi\cdot2=4\pi;~~~""\text{Answer:}~V_x=\dfrac{32}{15}\pi,~V_y=4\pi."